OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. 3x+4y=-2 2x-3y=10

OpenStudy (anonymous):

3x + 4y = -2 and 2x - 3y = 10 Multiply by 3 & 4 9x + 12y = -6 8x - 12y = 40 ADD 17x = 34 x = 2 y = -2